### v1

@@
Abstract
Security vulnerabilities in software packages are a significant concern for developers and users alike. Patching these vulnerabilities in a timely manner is crucial to restoring the integrity and security of software systems. However, previous work has shown that vulnerability reports often lack proof-of-concept (PoC) exploits, which are essential for fixing the vulnerability, testing patches, and avoiding regressions. Creating a PoC exploit is challenging because vulnerability reports are informal and often incomplete, and because it requires a detailed understanding of how inputs passed to potentially vulnerable APIs may reach security-relevant sinks. In this paper, we present PoCGen, a novel approach to autonomously generate and validate PoC exploits for vulnerabilities in npm packages. This is the first fully autonomous approach to use large language models (LLMs) in tandem with static and dynamic analysis techniques for PoC exploit generation. PoCGen leverages an LLM for understanding vulnerability reports, for generating candidate PoC exploits, and for validating and refining them. Our approach successfully generates exploits for 77% of the vulnerabilities in the SecBench.js dataset and 39% in a new, more challenging dataset of 794 recent vulnerabilities. This success rate significantly outperforms a recent baseline (by 45 absolute percentage points), while imposing an average cost of \$0.02 per generated exploit.
Index Terms:
@@ I Introduction
Security vulnerabilities pose a major threat to software and users alike, with the number of reported vulnerabilities increasing each year. In 2024 alone, over 40,000 CVEs were disclosed, an increase of 38% over the previous year [1]. As software ecosystems become more complex and interdependent, mitigating vulnerabilities becomes increasingly challenging. This holds particularly for Node.js and its package manager, npm, which form the backbone of modern JavaScript and TypeScript development. With millions of packages and a dense network of dependencies, the npm ecosystem is susceptible to a wide range of security risks [2], including transitive vulnerabilities, where a single vulnerable package can propagate security risks across thousands of applications [3].
@@
When a vulnerability is discovered, it is typically reported to the developers of the affected package, who are then expected to create a patch to fix the issue. Once the vulnerable software is fixed, or some time has passed from the initial vulnerability report, the vulnerability report is published as a Common Vulnerabilities and Exposures (CVE) entry. The process of fixing vulnerabilities is often facilitated by a proof-of-concept (PoC) exploit, which demonstrates how the vulnerability can be exploited in practice. Moreover, PoC exploits are useful for testing the patch and preventing regressions in the future. However, many vulnerability reports lack a PoC exploit [4], and even many CVE reports do not have any PoC exploit. For example, in the SecBench.js [5] dataset, only 179 out of 560 CVEs contain a PoC exploit in the report. Furthermore, the publicly available exploits are not reliable, and in some cases are malicious themselves [6].
Figure 1: CVE-2024-57063 report with no PoC exploit.
{minted}
[breaksymbol=, breaklines=true]markdown A prototype pollution in the lib function of php-date-formatter v1.3.6 allows attackers to cause a Denial of Service (DoS) via supplying a crafted payload.
@@
As a real-world example, the vulnerability CVE-2024-57063 in the “php-date-formatter” package shown in Fig. 1 does not contain a PoC exploit in the report. PoC exploits are useful for preventing regressions and identifying partial or ineffective fixes that fail to resolve the underlying vulnerability. The process of creating PoC exploits is often time-consuming and requires a deep understanding of the codebase, the vulnerability, and the underlying technology [5]. Particularly in the case of Fig. 1, the vulnerability description does not mention which function is vulnerable, and does not provide any information about the input that triggers the vulnerability.
With LLMs’ abilities in understanding and generating source code in addition to natural language, they have shown great promise in various software engineering tasks, including test generation, and program repair [7, 8, 9]. Hence, LLMs are good candidates to address the challenges in generating PoC exploits from vulnerability reports, specifically when the reports are vague. Moreover, with their understanding of vulnerability types, they can generate payloads for exploits in a more targeted manner, compared to traditional approaches that rely on symbolic execution or fuzzing [10, 11].
Recently, Marques et al. [12] proposed a method to find four vulnerability types in Node.js packages and generate exploits for them in a tool called Explode.js. Their approach uses taint analysis, a set of exploit templates, and symbolic execution to trigger vulnerabilities and output the exploits. Their work addresses the problem of generating PoC exploits for vulnerable npm packages. However, their approach does not use the vulnerability report and tries to find the vulnerabilities with taint analysis.
Since the recall of their approach on average is 58.2% and at most 66.2% for command injection vulnerabilities, there are vulnerabilities detected by developers or security researchers that are not detected by their approach. Therefore, generating PoC exploits from natural language description is of value to the community. Moreover, the exploit rate of their approach is on average 46.1% and at most 51.8% for prototype pollution vulnerabilities. Even for the vulnerabilities that are detected by their approach, on average 22% cannot be exploited by Explode.js.
@@
This paper aims to address these limitations by utilizing LLMs and program analysis techniques to automate the generation of PoC exploits from vulnerability reports in npm packages. Our proposed approach, PoCGen, takes as input an informal description of a vulnerability, as found in CVE reports, and automatically generates an executable exploit using a combination of LLM prompting and static and dynamic analysis. PoCGen consists of four, iteratively executed components: (i) understanding the vulnerability and extracting source-level information, (ii) generating the exploit, (iii) validating the exploit, and (iv) refining the prompt. Component (i) uses dynamic analysis to explore the package’s exported functions and prompts the LLM with the given vulnerability report to identify the vulnerable function. Component (ii) generates a candidate exploit using an LLM, which receives a prompt that contains a taint path from the vulnerable function to the exploit target. Component (iii) executes and validates the candidate exploit against a test oracle. If the candidate exploit is not valid, PoCGen uses component (iv) and refines the prompt using a set of refiners that provide static or dynamic information to component (ii) where the LLM attempts again to generate a valid exploit. This process continues until either a valid exploit is generated or a predefined budget is exhausted.
[…]
Figure 2: PoCGen-generated PoC exploit for CVE-2024-57063.
{minted}
javascript async function exploit() const DateFormatter=require(”php-date-formatter”); const maliciousJson=’”__proto__”: ”exploited”: true’; const maliciousOptions=JSON.parse(maliciousJson); const result=new DateFormatter(maliciousOptions); await exploit();
We run PoCGen on SecBench.js, a benchmark of currently 600 vulnerable npm packages with path traversal, prototype pollution, command injection, code injection, and ReDoS vulnerabilities. Since the SecBench.js benchmark only contains vulnerabilities up to 2022, we also evaluate PoCGen against a new dataset, that we extract from GitHub Advisory Database [^2] https://github.com/advisories and Snyk Vulnerability Database [^3] https://security.snyk.io/, which includes more recent vulnerabilities. Our results show that PoCGen successfully generates exploits for 432 out of 560 vulnerabilities in the SecBench.js dataset and 312 out of 794 in our new dataset. Our approach outperforms the state of the art by 45% on SecBench.js, while incurring the cost of only \$0.02 per vulnerability.
@@
By using PoCGen, developers of npm packages can generate PoC exploits for vulnerability reports they receive to help them understand the vulnerability and how to address it. They can also use the PoC exploits to test their patches and even add them to their test suites as regression tests. Moreover, security researchers can automate reporting vulnerabilities to downstream packages, by automating the PoC exploit generation process.
@@
  • A novel approach to autonomously generate and validate PoC exploits for vulnerabilities in npm packages using LLMs.
  • Empirical evidence of the effectiveness of PoCGen on the SecBench.js dataset and more recent vulnerabilities.
  • A new dataset of 794 vulnerabilities in npm packages, which includes more recent vulnerabilities and 312 PoC exploits generated by PoCGen.
  • Insights into the influence of vulnerability type on the success of exploit generation.
II Approach
Figure 3: Overview of PoCGen.
Figure 3 provides an overview of PoCGen. In summary, PoCGen consists of four main components: (i) vulnerability information extraction, (ii) exploit generation, (iii) validation, (iv) and prompt refinement, which are described in detail in the following sections.
PoCGen takes as input a vulnerability report, which is an informal description of the vulnerability. Such reports originate from several sources, such as vulnerability databases (e.g., the CVE database, GitHub Security Advisories, and Snyk), bug and issue trackers, or security mailing lists. The natural language description of the vulnerability in these reports is often vague and does not provide enough information to directly generate a PoC exploit. Therefore, in the first phase of our approach, PoCGen extracts some additional information from the vulnerability report and the codebase.
PoCGen extracts the vulnerability type, the likely vulnerable function, taint paths to vulnerable sinks, and usage examples. It then compiles them into a prompt for the LLM to generate a PoC exploit.
Once the LLM generates a candidate exploit, PoCGen executes and validates the candidate exploit using a set of runtime checkers, specific to each vulnerability type. For example, for command injection vulnerabilities, the validation checks whether a specific command is executed after running the candidate exploit, while for prototype pollution vulnerabilities, it checks whether a specific property is added to the global object.
Since the exploits are generated by LLMs, it is possible that the exploit does not use the vulnerable function to trigger the vulnerability, and calls some built-in functions that result in the same outcome. For example, in case of command injection, the call to exec(<some command>) can also result in the same outcome, but it does not invoke the vulnerable function. PoCGen checks the execution stack to detect such cases, and discards them as invalid exploits. Finally, to remove any invalid exploit that passed the validation, PoCGen checks whether the exploit actually triggers the vulnerability, by prompting the LLM with the vulnerability report and the generated PoC exploit.
If the candidate exploit is not valid, PoCGen refines the prompt using a set of refiners that provide static or dynamic information to the LLM to try to generate a valid exploit again. This process continues until either a valid exploit is generated or the budget is exhausted. If all validation checks pass and the LLM determines that the exploit correctly triggers the vulnerability, our approach considers it a valid PoC exploit.
II-A Vulnerability Information Extraction
@@ II-A1 Vulnerability Type
II-A2 Vulnerable Function
In our approach and throughout this paper, we refer to the function that is accessible to an attacker and is the entry point for the exploit as the vulnerable function. Finding the vulnerable function is helpful for generating the PoC exploit, as it provides information about the input types and possible values, and potentially the vulnerable execution path of the function. To identify the vulnerable function, we first load the package and dynamically extract all the functions that it exports. We then prompt the LLM with the vulnerability report and ask it to identify the vulnerable function from the extracted functions. All prompts and LLM responses are available in our supplementary material. Since the vulnerability report can be vague, and multiple functions can be candidates for the vulnerable function, we prompt the LLM to rank the functions based on their likelihood of being the vulnerable function.
II-A3 Taint Path Extraction
To generate a successful PoC exploit, it is crucial to understand how the input to the vulnerable function flows through the code and reaches the sensitive operations, commonly referred to as vulnerable sinks. Taint analysis is a technique that tracks the flow of data through the code by marking a certain input as tainted and propagating this taint through the code. We use the static taint analysis provided by CodeQL [^4] https://codeql.github.com/.
@@ II-A3 Taint Path Extraction
For each vulnerability type, we use the vulnerable sinks and taint propagation rules specified in the JavaScript security library of CodeQL [^5] The module semmle.javascript.security.
@@
The output of this phase is a sequence of code snippets interleaved with natural language explanations of how the code propagates the taint, shown in Section II-A3, with multiple sections if the taint path spans multiple files. In each section, there is a header specifying the file, followed by the taint path and its additional context from that file. The taint path lines are also marked by a comment at the end of the line.
@@
Figure 4: Example of a taint path extracted by PoCGen.
{minted}
@@
markdown Vulnerable method ‘import‘ of class ‘Environment‘ located in djv/lib/djv.js: “‘js import(config) // tainted: ”config” const item=JSON.parse(config) // tainted: ”config” let restoreData=item // tainted: ”item” if (item.name & item.fn item.schema) restoreData= [item.name]: item, Object.keys(restoreData).forEach((key)=¿ // tainted: ”restoreData” const name, schema, fn: source=restoreData[key] // tainted: ”restoreData” const fn=restore(source, schema, this.options) // tainted: ”source” this.resolved[name]= name, schema, “‘ Call to ‘restore‘: “‘js function restore(source, schema, inner=) // tainted: ”source” const tpl=new Function(”schema”, source)(schema) // tainted: ”source” if (!inner) “‘
II-A4 Usage Snippets

### v3

@@
Abstract.
Security vulnerabilities in software packages are a significant concern for developers and users alike. Patching these vulnerabilities in a timely manner is crucial to restoring the integrity and security of software systems. However, previous work has shown that vulnerability reports often lack proof-of-concept (PoC) exploits, which are essential for fixing the vulnerability, testing patches, and avoiding regressions. Creating a PoC exploit is challenging because vulnerability reports are informal and often incomplete, and because it requires a detailed understanding of how inputs passed to potentially vulnerable APIs may reach security-relevant sinks. In this paper, we present PoCGen, a novel approach to autonomously generate and validate PoC exploits for vulnerabilities in npm packages. The approach is the first to address this task by combining the complementary strengths of large language models (LLMs), e.g., to understand informal vulnerability reports, with static analysis, e.g., to identify taint paths, and dynamic analysis, e.g., to validate generated exploits. PoCGen successfully generates exploits for 77% of the vulnerabilities in the SecBench.js dataset. This success rate significantly outperforms a recent baseline (by 45 absolute percentage points), while imposing an average cost of only \$0.02 per generated exploit. Moreover, PoCGen generates six successful exploits for recent real-world vulnerabilities, five of which are now included in their respective vulnerability reports.
@@ 1. Introduction
Security vulnerabilities pose a major threat to software and users alike, with the number of reported vulnerabilities increasing each year. In 2024 alone, over 40,000 CVEs were disclosed, an increase of 38% over the previous year (CVE, [n. d.]). As software ecosystems become more complex and interdependent, mitigating vulnerabilities becomes increasingly challenging. This holds particularly for Node.js and its package manager, npm, which form the backbone of the JavaScript and TypeScript ecosystem. With millions of packages and a dense network of dependencies, the npm ecosystem is susceptible to a wide range of security risks (Zimmermann et al., 2019), including transitive vulnerabilities, where a single vulnerable package can propagate security risks across thousands of applications (Zahan et al., 2022).
@@
When a vulnerability is discovered, it is typically reported to the developers of the affected package, who are then expected to create a patch to fix the issue. Once the vulnerable software is fixed, or some time has passed from the initial vulnerability report, the vulnerability report is published as a Common Vulnerabilities and Exposures (CVE) entry. The process of fixing vulnerabilities is often facilitated by a proof-of-concept (PoC) exploit, which demonstrates how the vulnerability can be exploited in practice. Moreover, PoC exploits are useful for testing the patch and preventing regressions in the future. However, many vulnerability reports lack a PoC exploit (Householder et al., 2020), and even many CVE reports do not have any PoC exploit. For example, in the SecBench.js (Bhuiyan et al., 2023) dataset, only 179 out of 560 CVEs contain a PoC exploit in the report. Furthermore, the publicly available exploits are not reliable, and in some cases are malicious themselves (Yadmani et al., 2023).
Figure 1. CVE-2024-36751 report with no PoC exploit in the report.
An issue in parse-uri v1.0.9 allows attackers to cause a Regular expression Denial of Service (ReDoS) via a crafted URL.
@@
As a real-world example, consider the vulnerability CVE-2024-36751 in the “parse-uri” package shown in Fig. 1. The vulnerability report informally describes the problem, but it does not contain an executable code example to reproduce it, i.e., there is no PoC exploit. PoC exploits are useful for preventing regressions and identifying partial or ineffective fixes that fail to resolve the underlying vulnerability. The process of creating PoC exploits is often time-consuming and requires a deep understanding of the codebase, the vulnerability, and the underlying technology (Bhuiyan et al., 2023). Particularly in the case of Fig. 1, the vulnerability description does not mention which function is vulnerable, and does not provide any information about the input that triggers the vulnerability.
One way to address the challenge of generating a PoC exploit that is missing in a given vulnerability report is to leverage the capabilities of large language models (LLMs). LLMs have demonstrated their effectiveness in various software engineering tasks, including code completion (Chen et al., 2021), test generation(Schäfer et al., 2024; Lemieux et al., 2023), program repair (Jiang et al., 2023; Jin et al., 2023; Xia and Zhang, 2024; Zhang et al., 2024a; Bouzenia et al., 2025), and repository setup (Bouzenia and Pradel, 2025). They are especially good candidates because of their ability to understand natural language and other informal, imprecise, and vague information provided in vulnerability reports. Moreover, with their understanding of different vulnerability types, LLMs can generate payloads for exploits in a targeted manner. However, LLMs alone may not be sufficient to generate successful PoC exploits, as they lack the necessary context about the codebase, such as implementation details of the vulnerable code, and have only a limited ability to reason about the behavior of the code.
Another way to generate PoC exploits is to deploy traditional testing and analysis techniques, such as fuzzing (Cassel et al., 2025) and symbolic execution (Marques et al., 2025). The currently most effective such technique is a recent approach called Explode.js (Marques et al., 2025), which uses taint analysis, a set of exploit templates, and symbolic execution to produce exploits that trigger vulnerabilities in npm packages. Explode.js has been shown to generate PoC exploits for 182 out of 560 vulnerabilities (32%) in the SecBench.js dataset, which is helpful, but still leaves many vulnerabilities without a PoC exploit. We attribute this gap to the fact that Explode.js uses traditional program testing and analysis techniques only, but leverages neither the information provided in vulnerability reports nor the capabilities of LLMs.
@@
This paper addresses these limitations by presenting PoCGen, a novel approach for generating PoC exploits from informal vulnerability reports. PoCGen is the first to address this problem by combining LLMs with static and dynamic program analysis. The approach takes as input an informal description of a vulnerability, as found in CVE reports, as well as the vulnerable code base, and automatically generates an executable JavaScript file that demonstrates how to exploit the vulnerability. PoCGen consists of four, iteratively executed components: (i) understanding the vulnerability and extracting source-level information, (ii) generating a candidate exploit, (iii) validating the exploit, and if necessary, (iv) refining the prompt to obtain an improved exploit. These components use a combination of LLM prompting and static and dynamic analysis techniques: Component (i) uses dynamic analysis to explore the package’s exported functions, prompts the LLM with the given vulnerability report to identify potentially vulnerable functions, and queries a static taint analysis to extract taint paths. Component (ii) generates a candidate exploit using an LLM. Component (iii) executes and validates the candidate exploit against a dynamic analysis-based test oracle. If the candidate exploit is not valid, PoCGen uses component (iv) and refines the prompt using a set of refiners that provide static or dynamic information to component (ii) where the LLM attempts again to generate a valid exploit. This process continues until either a valid exploit is generated or until exceeding the computational budget.
To generate a PoC exploit for the ReDoS vulnerability in Fig. 1, the main challenge is to construct a payload that, when passed to the vulnerable function, triggers a security-relevant action. The goal of ReDoS is to exploit the regular expression engine’s backtracking behavior, leading to excessive resource consumption. After analyzing the codebase for potential vulnerable functions, PoCGen identifies that there is a default exported function, and extracts some usage examples from the codebase as well. In its first attempt, PoCGen generates an exploit that passes a crafted input string designed to trigger the ReDoS vulnerability. However, this payload is invalid as it does not cause the expected backtracking behavior. Once PoCGen executes the initial candidate exploit, it realizes that the exploit does not work, and by reasoning about the vulnerability, the code, and runtime information, the LLM comes up with a new payload. After multiple refinements, PoCGen generates a new exploit that triggers the ReDoS vulnerability, as shown in Fig. 2.
Figure 2. PoCGen-generated PoC exploit for CVE-2024-36751.
async function exploit() {
const parseuri = require(”parse-uri”);
// This input is designed to cause excessive backtracking in the regex
const craftedInput = http://example.com/’ + a’.repeat(30000) + ’?key=value’;
const result = await parseuri(craftedInput);
await exploit();
We evaluate PoCGen on SecBench.js, a benchmark of vulnerable npm packages with path traversal, prototype pollution, command injection, code injection, and ReDoS vulnerabilities. Our results show that PoCGen successfully generates exploits for 432 out of 560 vulnerabilities (77%), outperforming the previous state of the art (Marques et al., 2025) by 45 absolute percentage points. When measuring the costs due to LLM usage, we find that PoCGen incurs an average cost of only \$0.02 per vulnerability, which is reasonably low for adoption in practice. To further validate the practical usefulness of PoCGen, we use it to generate PoC exploits for six recent real-world vulnerabilities that previously lacked a PoC exploit in their reports. Five of the six generated exploits have been accepted and added to their respective vulnerability reports, and the remaining one is still pending.
@@
We envision the approach to be useful for developers and security researchers. By using PoCGen, developers of npm packages can generate PoC exploits for vulnerability reports they receive to help them understand the vulnerability and how to address it. They can also use the PoC exploits to test their patches and even add them to their test suites as regression tests. Moreover, security researchers can automate reporting vulnerabilities to downstream packages, by automating the PoC exploit generation process.
@@
  • A novel approach to autonomously generate and validate PoC exploits for vulnerabilities in npm packages by combining LLMs with static and dynamic analysis.
  • Empirical evidence of the effectiveness of PoCGen on 560 real-world vulnerabilities, covering five common vulnerability types, as well as its utility for augmenting vulnerability reports that were previously missing PoC exploits.
  • We share our code and data to foster future research (Section 6).
2. Approach
2.1. Overview
Figure 3. Overview of PoCGen.
Figure 3 provides an overview of PoCGen. The approach consists of four main components, which combine LLM prompting, static analysis, and dynamic analysis. PoCGen takes as input a vulnerability report, which is an informal description of the vulnerability, and the codebase of the vulnerable package. Vulnerability reports originate from several sources, such as vulnerability databases (e.g., the CVE database, GitHub Security Advisories, and Snyk), bug and issue trackers, or security mailing lists.
The first step of PoCGen (Section 2.2) is to extract information about the vulnerability and the codebase. This step is necessary because the natural language description in a vulnerability report is often vague and does not provide enough information to directly generate a PoC exploit. The approach extracts the vulnerability type, the likely vulnerable function, taint paths to vulnerable sinks, and usage examples.
In the second step (Section 2.3), PoCGen uses the extracted information to generate a candidate exploit. To this end, the approach compiles all the available information into a prompt that asks the LLM to generate a PoC exploit.
Once the LLM generates a candidate exploit, the third step of PoCGen (Section 2.4) executes and validates the candidate exploit. The approach uses multiple validation mechanisms, including a a set of runtime checkers that are specific to each vulnerability type. For example, for command injection vulnerabilities, the validation checks whether a specific command is executed after running the candidate exploit, while for prototype pollution vulnerabilities, it checks whether a specific property is added to the global object.
In case the candidate exploit is successfully validated, PoCGen returns it to the user and the approach terminates. Otherwise, the fourth step of the approach (Section 2.5) refines the exploit generation prompt by adding additional context based on static analysis and the runtime information obtained during exploit generation. This process continues until either finding a valid exploit or exhausting the computational budget.
2.2. Vulnerability Information Extraction
@@ 2.2.1. Vulnerability Type
2.2.2. Vulnerable Function
In our approach and throughout this paper, we refer to the function that is accessible to an attacker and is the entry point for the exploit as the vulnerable function. Finding the vulnerable function is helpful for generating the PoC exploit, as it provides information about the input types and possible values, and potentially the vulnerable execution path of the function. To identify the vulnerable function, we first load the package and dynamically extract all the functions that it exports. We then prompt the LLM with the vulnerability report and ask it to identify the vulnerable function from the extracted functions. Since the vulnerability report can be vague, and multiple functions can be candidates for the vulnerable function, we prompt the LLM to rank the functions based on their likelihood of being the vulnerable function. This ranked list of candidate functions is the input to the next step which is to extract taint paths.
2.2.3. Taint Path Extraction
To generate a successful PoC exploit, it is crucial to understand whether and how the input to the vulnerable function flows through the code and reaches the sensitive operations, commonly referred to as vulnerable sinks. Taint analysis is a technique that tracks the flow of data through the code by marking a certain input as tainted and propagating this taint through the code. Moreover, to narrow down the search space for the vulnerable function, PoCGen attempts to extract a taint path from each of the candidate functions identified in the previous step to a vulnerable sink. If no taint path is found for a candidate function, that function is not considered for exploit generation.
@@ 2.2.3. Taint Path Extraction
We use the static taint analysis provided by CodeQL [^1] https://codeql.github.com/. For each vulnerability type, we use the vulnerable sinks and taint propagation rules specified in the JavaScript security library of CodeQL [^2] The module semmle.javascript.security.
@@
The output of this phase is a sequence of code snippets interleaved with natural language explanations of how the code propagates the taint, shown in Section 2.2.3, with multiple sections if the taint path spans multiple files. In each section, there is a header specifying the file, followed by the taint path and its additional context from that file. The taint path lines are also marked by a comment at the end of the line.
The package djv in Section 2.2.3, which is a JSON schema validator, had a code injection vulnerability (CVE-2020-28464). The vulnerable function import takes a JSON string as input, parses it, and then calls the restore function with each of the attributes of the parsed object. In the restore function, a new function is dynamically created where the body of the function comes from the parsed JSON string. As shown in Section 2.2.3, a malicious input (e.g., JSON string containing arbitrary JavaScript code) to the import function, flows through to the restore function, and is used as the body of a dynamically created function, which is a code injection vulnerability.
@@
Figure 4. Example of a taint path extracted by PoCGen from the package djv with code injection vulnerability (CVE-2020-28464).
@@
Vulnerable method ‘import‘ of class ‘Environment‘ located in djv/lib/djv.js: “‘js import(config) // tainted: ”config” const item=JSON.parse(config) // tainted: ”config” let restoreData=item // tainted: ”item” if (item.name item.fn item.schema) restoreData= [item.name]: item, Object.keys(restoreData).forEach((key)=¿ // tainted: ”restoreData” const name, schema, fn: source=restoreData[key] // tainted: ”restoreData” const fn=restore(source, schema, this.options) // tainted: ”source” this.resolved[name]= name, schema, “‘ Call to ‘restore‘: “‘js function restore(source, schema, inner=) // tainted: ”source” const tpl=new Function(”schema”, source)(schema) // tainted: ”source” if (!inner) “‘
2.2.4. Usage Snippets